5 years ago a son was 1/3 the age of his father. 5 years from now he will be 1/2 the age of his father. How old are both now?
Answer by scott8148(6628)

let x="son now" and y="father now"

"5 years ago a son was 1/3 the age of his father" __ 3(x-5)=y-5 __ 3x-15=y-5

"5 years from now he will be 1/2 the age of his father" __ 2(x+5)=y+5 __ 2x+10=y+5

subtracting second equation from first __ x-25=-10

substitute the value of x into either equation to find y